Do you know how to solve that system? If not
post again asking how. The solutions are
A = 1/6, B = 0, C = 11/6, D = -1

Answer by ikleyn(52787)

.
Do you know how to solve that system? If not
post again asking how. The solutions are
A = 1/6, B = 0, C = 11/6, D = -1
~~~~~~~~~~~~~~


    G I B B E R I S H.



Hey, If you will post gibberish to this forum, I will contact with your supervisors.


Asking them to remove you from your position.